# Benchmarking Cross-Domain Audio-Visual Deception Detection

### v2

@@ 4.2 Implementation Details
Feature Extraction. Several widely-adopted audio and visual features were extracted by different tools. For visual features, OpenFace [36] was used to extract 35-dimensional AUs and 8-dimensional gaze features. Face frames were extracted and aligned by MTCNN [45], where we uniformly sampled 64 face frames for each video clip. Affect features were extracted by Emonet [37], where the feature included 5-class emotions, arousal, and valence. For audio features, Mel Spectrograms were extracted by OpenSmile toolkit [38]. Raw audio waveforms were also used in our experiments.
Protocols. Inspired by [39, 40], we treated each dataset as a domain. To evaluate the models’ cross-domain generalization capacity and alleviate domain information leakage, all the preprocessed data including original training and test data from each dataset was used for either training or testing. Note that the Box of Lies dataset was only used for testing as many samples were filtered out due to their unsatisfactory quality. The experiments were conducted on the single-to-single domain (e.g., R to B1 stands for training on Real-life Trial (R) and testing on Bag of Lies (B1)) and multi-to-single domain (e.g., R&M to B2 stands for training on Real-life Trial (R) and MU3D (M) and testing on Box of Lies (B2)).
Model Selection. Models for audio and visual modalities were selected to fit the data volume. For face frames, we adopted ResNet18 [35] and Gate Recurrent Unit (GRU) [46] models for facial feature extraction and temporal modeling, respectively. Two-layer multilayer perception (MLP) [47] models were used for AUs, gaze, and affect feature representation. Specifically, the first MLP Layer initially operates on the dimension that represents the number of features, with sizes following the sequence: “input dimension $\rightarrow$ 64 $\rightarrow$ 128 $\rightarrow$ 1. Afterward, it reshapes the tensor, and the MLP Layer processes on the feature dimension, where the sizes progress as “64 $\rightarrow$ 32 $\rightarrow$ 16 $\rightarrow$ 2. The MLP networks have identical structures, except for their first linear layers, as the input dimensions differ. For the audio-based Mel spectrogram, we used the ResNet18 [35] model for time-frequency feature representation. For audio waveforms, the Wave2Vec [48] model was applied for audio feature extraction.
@@
TABLE I: The results of intra-domain testing accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), and MU3D (M).
Modality & Inputs Method R to R B1 to B1 M to M
Visual (Face frames) ResNet18 70.10 57.78 55.15
Visual (AU+Gaze+Affect) MLP 62.87 55.84 55.23
Audio (Mel spectrogram) ResNet18 75.78 55.69 56.45
Visual (Face frames) concat 59.09 54.46 56.30
+ Visual (AU+Gaze+Affect) Atten-Mixer (Ours) 63.05 55.45 57.59
Visual (Face frames) concat 68.75 54.74 56.19
+ Audio (Mel spectrogram) Atten-Mixer (Ours) 71.00 56.44 57.14
Visual (AU+Gaze+Affect) concat 68.88 56.25 53.27
+ Audio (Mel spectrogram) Atten-Mixer (Ours) 69.30 56.78 55.23
Visual (Face frames) concat 69.33 58.43 55.24
+ Visual (AU+Gaze+Affect) Atten-Mixer(Ours) 70.52 59.41 56.88
+ Audio (Mel spectrogram)
TABLE II: The results of single-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M).
Modality & Inputs Method R to B1 R to B2 R to M B1 to R B1 to B2 B1 to M M to R M to B1 M to B2 Avg
Visual (AU) LSTM 48.11 - - 61.21 - - - - - -
Visual (Face frames) ResNet18 52.00 61.39 51.25 50.93 57.43 50.62 57.94 51.69 57.43 54.52
Visual (Face frames) ResNet18+GRU 53.54 63.37 52.81 57.41 59.41 51.56 46.73 52.92 55.45 54.80
Visual (AU+Gaze) MLP 50.77 65.35 56.87 58.88 58.42 50.94 46.73 51.69 53.47 54.79
Visual (Affect) MLP 50.46 58.42 50.31 50.47 51.49 52.19 66.36 51.08 60.40 54.58
Visual (AU+Gaze+Affect) MLP 54.46 59.41 54.37 50.47 57.43 54.69 60.75 51.69 55.45 55.41
Audio (Mel spectrogram) ResNet18 46.77 53.47 52.19 50.47 66.34 50.62 54.21 51.38 55.45 53.43
Audio (Waveform) Wave2Vec 51.08 48.51 50.94 46.73 58.42 50.00 63.55 56.31 56.44 53.55
TABLE III: The results of multi-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M), for different generalization strategies.
Modality & Inputs Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 Avg
@@
Visual (Face frames) ResNet18 53.85 49.50 49.50 50.94 44.86 60.40 44.55 50.51
Visual (Face frames) ResNet18+GRU 52.62 54.46 51.49 51.88 53.27 59.41 44.55 52.53
Visual (AU+Gaze) MLP 53.54 47.52 48.51 50.94 52.34 53.47 56.44 51.82
Visual (Affect) MLP 50.15 54.46 55.45 52.19 53.27 57.43 62.38 55.04
Visual (AU+Gaze+Affect) MLP 50.46 52.48 61.39 51.25 51.4 60.4 63.37 55.82
Audio (Mel spectrogram) ResNet18 48.92 45.54 53.47 53.12 43.93 62.38 50.5 51.12
Audio (Waveform) Wave2Vec 52.92 55.45 44.55 51.25 69.16 42.57 46.53 51.78
@@
Visual (Face frames) ResNet18 50.15 45.54 56.44 51.56 50.47 54.46 65.85 53.50
Visual (Face frames) ResNet18+GRU 55.38 52.48 60.40 50.00 50.47 60.40 64.62 56.25
Visual (AU+Gaze) MLP 55.45 47.52 53.47 51.25 54.21 57.43 60.40 54.25
Visual (Affect) MLP 51.08 56.44 61.39 52.19 52.34 58.42 53.47 55.05
Visual (AU+Gaze+Affect) MLP 51.38 58.42 63.37 50.31 53.27 52.48 60.40 55.66
Audio (Mel spectrogram) ResNet18 50.15 60.40 53.47 50.31 58.88 51.49 47.52 53.17
Audio (Waveform) Wave2Vec 52.92 55.45 44.55 50.62 64.49 58.42 48.51 53.57
@@
Visual (Face frames) ResNet18 52.00 53.47 56.44 50.00 59.81 41.58 55.45 52.68
Visual (Face frames) ResNet18+GRU 54.46 41.58 66.34 50.62 51.40 56.44 60.40 54.46
Visual (AU+Gaze) MLP 51.08 43.56 55.45 53.75 57.01 53.47 54.46 52.68
Visual (Affect) MLP 55.69 57.43 57.43 51.56 52.34 49.50 61.39 55.05
Visual (AU+Gaze+Affect) MLP 50.15 56.44 58.42 50.00 57.94 60.40 63.37 56.67
Audio (Mel spectrogram) ResNet18 52.31 50.50 58.42 49.38 53.27 56.44 59.41 54.24
Audio (Waveform) Wave2Vec 56.00 47.52 44.55 53.12 67.29 57.43 58.42 54.90
TABLE IV: The fusion results of single-to-single cross-domain generalization accuracy (%).
Modality & Inputs Fusion Postion Fusion Method R to B1 R to B2 R to M B1 to R B1 to B2 B1 to M M to R M to B1 M to B2 Avg
Score-level Average 53.23 41.58 51.88 64.49 62.38 50.62 57.94 53.23 62.38 55.30
Concat 51.08 55.45 51.88 54.21 58.42 51.25 57.01 50.15 61.39 54.54
SE-Concat 53.85 60.40 51.25 55.14 58.42 50.62 56.07 51.38 65.35 55.83
Cross-Atten 55.38 61.39 52.19 51.40 60.40 51.25 55.14 56.31 60.40 55.98
MLP-Mixer 55.08 48.51 53.44 56.07 58.42 53.75 55.14 59.08 59.41 55.43
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 56.92 59.41 57.94 63.37 53.75 53.75 60.75 56.00 61.39 58.14
Score-level Average 53.23 49.50 51.88 50.47 59.41 53.75 65.42 56.31 49.50 54.39
Concat 50.77 53.47 50.47 62.38 51.56 51.88 54.21 52.62 58.42 53.98
SE-Concat 50.15 44.55 51.40 61.39 53.12 52.50 65.42 56.31 66.34 55.69
Cross-Atten 54.46 51.49 55.14 58.42 51.25 52.19 63.55 56.62 66.34 55.95
MLP-Mixer 52.31 55.45 57.94 63.37 53.12 51.25 64.49 57.85 62.38 57.57
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 57.54 55.45 56.07 61.39 50.94 53.12 67.29 57.85 64.36 58.22
Score-level Average 49.85 58.42 54.06 45.79 60.40 50.62 49.53 56.00 63.37 54.23
Concat 49.54 53.47 47.66 61.39 53.44 51.88 57.01 50.15 58.42 53.66
SE-Concat 50.15 48.51 55.14 60.40 53.12 50.00 57.01 49.85 63.37 54.17
Cross-Atten 53.23 44.55 57.94 55.45 54.69 54.06 63.55 51.69 64.36 55.50
MLP-Mixer 49.54 57.43 50.47 63.37 54.06 53.12 59.81 55.08 69.31 56.91
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 53.52 54.46 57.94 61.39 53.12 51.56 69.16 58.15 64.36 58.18
Score-level Average 52.00 58.42 51.88 53.27 58.42 50.94 57.01 53.54 61.39 55.20
Concat 53.23 58.42 55.14 61.39 52.50 52.19 56.07 54.15 60.40 55.94
SE-Concat 51.38 58.42 51.40 62.38 52.81 50.94 59.81 54.77 52.48 54.93
Cross-Atten 51.08 48.51 55.14 60.40 53.12 53.44 60.75 56.31 60.40 55.46
MLP-Mixer 55.69 46.53 44.86 63.37 51.56 50.94 64.49 56.00 60.40 54.87
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.08 60.40 57.01 64.36 53.44 51.25 67.29 56.00 62.38 58.58
@@ 4.3 Cross-domain Testing with Unimodal Features
Intra-Domain Testing. Before performing cross-domain testing, we show the results of intra-domain testing on Real-life Trial (R), Bag of Lies (B1), and MU3D (M) in Table I. It is important to note that we did not use the Box of Lies (B2) dataset for training due to concerns about its quality. Specifically, we performed 5-fold intra-domain testing and reported the average accuracies. ResNet18 and MLP were used as baseline models, as they are widely used as encoders in audio-visual deception detection. As expected, the performance of intra-domain testing varies across datasets, reflecting the differences in their levels of difficulty.
Single-to-Single Domain. Specifically, the models were trained on one dataset from one domain and tested on the other dataset from another domain. The experiments were conducted on the four public datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M). As shown in Tabel II, for visual modalities, we extracted the most adopted visual features including face frames, and behavior features such as AUs, gaze and affect. For audio modality, Mel spectrogram and waveform were extracted. We also applied several different backbone networks as audio and visual encoders. We can observe that R and B1 datasets generalized the best on B2, and M generalized the best on R. Note that the B2 dataset was not adopted as a source domain dataset as the original dataset had too much noise and we cleaned it only for testing. On average, we can observe that the best result was achieved by using visual (AU+gaze+affect) features.
Multi-to-Single Domain. Here we fully evaluate the performance of multi-to-single cross-domain generalization using unimodal features. We conducted experiments for different domain sampling strategies. As shown in Table III, for domain-simultaneous training, the best generalization performance is achieved by training on the Bag of Lies and MU3D datasets and testing on the Real-life Trial dataset (69.16%), which was also the case for domain-by-domain training strategy (67.29%). For the domain-alternating strategy, the best result was observed when transferring to the Box of Lies dataset using the rest three datasets for training (65.85%). The results showed that the best generalization performances were obtained when transferring to the real-world dataset and the gameshow dataset by training on the lab-based datasets. This was because the lab-based datasets (Bag of Lies and MU3D) are relatively clean compared to the real-world dataset (Real-life Trial) and the gameshow dataset (Box of Lies). However, in the opposite case, the generalization performance degraded, for example, R&M to B1 and R&B1 to M. Different domain sampling strategies reached their best average performance on different input features and backbone networks. To be specific, for both domain-simultaneous and domain-by-domain strategies, models trained on visual (AU+gaze+affect) features reached their highest accuracies, which were 55.82% and 56.67%, respectively. Using the domain-alternating strategy, the best accuracy of 56.25% was achieved by training on visual (face frames) features. We can observe that models trained on visual modalities outperformed those trained on audio modalities across all the generalization strategies. This may be due to the rich deceptive cues captured by visual modalities in the publicly available datasets.
Discussion: (1) Wave2Vec outperforms MFCC in audio feature extraction due to its ability to capture richer, more meaningful representations, as demonstrated across various sampling methods. Additionally, larger and more diverse datasets like B1 and M enhance generalization, leading to improved performance when models trained on these datasets are tested on smaller or less challenging datasets like R. The R dataset, composed of courtroom trial videos, tends to yield better results, while B2, a game show dataset, often produces unsatisfactory outcomes due to its lower video quality and data imbalance. Overall, visual features perform better than audio features and models using Wave2Vec features achieve better results compared to those trained on MFCC.
(2) While domain-simultaneous sampling is typically expected to enhance generalization by promoting domain-invariant learning, our average results show that domain-by-domain sampling slightly outperforms it in these audio-visual deception detection tasks. This may be due to its ability to capture domain-specific variations, which are beneficial in this context. However, the effectiveness of each sampling strategy depends on the features being used; for example, Wave2Vec features perform better with domain-simultaneous sampling, while MFCC features benefit more from domain-by-domain sampling. This suggests that the optimal sampling strategy may vary depending on the feature set’s predictive power and the specific dataset characteristics.
(3) Visual features, particularly behavioral features like AUs, gaze, and affective cues, outperform audio features in most of our experiments. Behavioral features are more predictive in the majority of testing pairs, as seen in Table II, and achieved the best average performance across sampling strategies in Table III. Their rich information content, capturing subtle signs of deception, may explain their superior performance over the face or audio features. Due to the lower dimension compared with face and audio features, behavioral features make them more effective when training on small datasets for detecting deception.

### v3

@@ 4.2 Implementation Details
Feature Extraction. Several widely-adopted audio and visual features were extracted by different tools. For visual features, OpenFace [39] was used to extract 35-dimensional AUs and 8-dimensional gaze features. Face frames were extracted and aligned by MTCNN [48], where we uniformly sampled 64 face frames for each video clip. Affect features were extracted by Emonet [40], where the feature included 5-class emotions, arousal, and valence. For audio features, Mel Spectrograms and acoustic features were extracted by the OpenSmile toolkit [41]. Raw audio waveforms were also used in our experiments.
Protocols. Inspired by [42, 43], we treated each dataset as a domain. To evaluate the models’ cross-domain generalization capacity and alleviate domain information leakage, all the preprocessed data including original training and test data from each dataset was used for either training or testing. Note that the Box of Lies dataset was only used for testing, as many samples were filtered out due to their unsatisfactory quality. Because of shifting camera angles and background audience noise, the speaker’s face and speech may not be captured clearly in the video. The experiments were conducted on the single-to-single domain (e.g., R to B1 stands for training on Real-life Trial (R) and testing on Bag of Lies (B1)) and multi-to-single domain (e.g., R&M to B2 stands for training on Real-life Trial (R) and MU3D (M) and testing on Box of Lies (B2)).
Model Selection. Different audio and visual features are used for benchmarking, including gaze, AUs, affect, facial landmark (2D, 68 points, using dlib package), acoustic and prosodic features (88-dimensional features extracted from the OpenSmile toolkit), Mel spectrograms, and audio waveforms. We also benchmark and re-implement the methods that are widely adopted by domain-specific deception detection task [20], such as ResNet18, MLP, SVM (Linear) [19, 25], KNN [24], FFCSN [7], LSTM [19], etc.
Models for audio and visual modalities were selected to fit the data volume. For face frames, we adopted ResNet18 [38], and Gate Recurrent Unit (GRU) [49], SVM [19, 25], and KNN [24] models for facial feature extraction and temporal modeling, respectively. Two-layer multilayer perception (MLP) [50] or LSTM [19] models were used for AUs, gaze, and affect feature representation. Specifically, the first MLP Layer initially operates on the dimension that represents the number of features, with sizes following the sequence: “input dimension $\rightarrow$ 64 $\rightarrow$ 128 $\rightarrow$ 1. Afterward, it reshapes the tensor, and the MLP Layer processes on the feature dimension, where the sizes progress as “64 $\rightarrow$ 32 $\rightarrow$ 16 $\rightarrow$ 2. The MLP networks have identical structures, except for their first linear layers, as the input dimensions differ. For acoustic features, MLP is used. For the audio-based Mel spectrogram, we used the ResNet18 [38], KNN, and SVM models for time-frequency feature representation and classification. For audio waveforms, the Wave2Vec [51] model was applied for audio feature extraction.
@@
TABLE IV: The results of multi-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M), for different generalization strategies.
@@
Modality & Inputs Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 R&M to D R&M to E R&B1 to D R&B1 to E
Visual (Face frames) ResNet18 53.85 49.50 49.50 50.94 44.86 60.40 44.55 52.43 50.88 54.82 49.92
Visual (Face frames) ResNet18+GRU 52.62 54.46 51.49 51.88 53.27 59.41 44.55 53.06 51.88 53.98 50.33
Visual (Face frames) ResNet18+SVM 52.98 54.87 52.08 52.66 54.12 59.39 45.38 53.56 52.39 54.08 51.24
Visual (AU+Gaze) MLP 53.54 47.52 48.51 50.94 52.34 53.47 56.44 50.48 51.28 53.35 50.22
Visual (Affect) MLP 50.15 54.46 55.45 52.19 53.27 57.43 62.38 52.08 52.33 53.24 51.02
Visual (AU+Gaze+Affect) MLP 50.46 52.48 61.39 51.25 51.4 60.40 63.37 52.92 53.45 53.99 52.27
Audio (Mel spectrogram) ResNet18 48.92 45.54 53.47 53.12 43.93 62.38 50.50 51.06 50.88 52.25 51.43
Audio (Mel spectrogram) ResNet18+SVM 49.88 47.99 52.64 52.96 46.98 59.88 51.28 52.04 51.36 53.82 52.77
Audio (Waveform) Wave2Vec 52.92 55.45 44.55 51.25 69.16 42.57 46.53 53.26 52.18 53.69 53.06
Modality & Inputs Method D&E to B1 D&E to B2 D&E to M D&E to R R&B1&M to D R&B1&M to E R&D&E to B1 R&D&E to M R&M&D&E to B2 R&B1&M&D&E to B2 Average
Visual (Face frames) ResNet18 50.21 52.39 51.09 51.39 49.55 50.03 51.88 50.58 51.84 50.79 51.02
Visual (Face frames) ResNet18+GRU 51.25 52.98 52.92 52.24 50.26 51.26 52.26 51.23 52.06 51.26 52.13
Visual (Face frames) ResNet18+SVM 52.56 53.26 52.36 53.62 51.22 51.48 53.18 51.46 52.47 51.63 52.67
Visual (AU+Gaze) MLP 50.48 52.48 52.02 51.98 50.23 51.06 51.66 50.23 51.02 50.16 51.40
Visual (Affect) MLP 51.35 53.45 52.88 53.26 52.45 52.48 52.87 51.45 52.38 51.26 53.23
Visual (AU+Gaze+Affect) MLP 52.48 54.28 54.33 54.18 52.35 53.18 61.46 53.29 54.39 53.28 54.60
Audio (Mel spectrogram) ResNet18 51.33 51.47 52.02 51.05 50.25 50.49 51.06 51.38 50.29 51.33 51.15
Audio (Mel spectrogram) ResNet18+SVM 52.55 52.39 52.78 52.49 51.62 51.63 52.68 52.58 51.64 52.68 52.13
Audio (Waveform) Wave2Vec 53.09 53.69 53.85 52.63 52.04 51.38 52.04 52.44 51.83 52.82 52.40
@@
Modality & Inputs Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 R&M to D R&M to E R&B1 to D R&B1 to E
Visual (Face frames) ResNet18 50.15 45.54 56.44 51.56 50.47 54.46 65.85 52.28 51.34 54.65 50.32
Visual (Face frames) ResNet18+GRU 55.38 52.48 60.40 50.00 50.47 60.40 64.62 53.66 52.26 54.03 51.29
Visual (Face frames) ResNet18+SVM 56.21 52.48 61.23 51.05 51.28 59.39 64.28 53.54 53.65 54.88 52.43
Visual (AU+Gaze) MLP 55.45 47.52 53.47 51.25 54.21 57.43 60.40 51.29 52.26 53.46 50.33
Visual (Affect) MLP 51.08 56.44 61.39 52.19 52.34 58.42 53.47 52.45 52.56 53.83 50.98
Visual (AU+Gaze+Affect) MLP 51.38 58.42 63.37 50.31 53.27 52.48 60.40 53.76 53.58 54.72 51.26
Audio (Mel spectrogram) ResNet18 50.15 60.40 53.47 50.31 58.88 51.49 47.52 52.02 51.63 54.22 50.36
Audio (Mel spectrogram) ResNet18+SVM 51.53 59.62 53.02 51.26 60.38 52.56 47.33 51.76 51.38 53.55 51.82
Audio (Waveform) Wave2Vec 52.92 55.45 44.55 50.62 64.49 58.42 48.51 53.26 52.35 53.29 52.47
Modality & Inputs Method D&E to B1 D&E to B2 D&E to M D&E to R R&B1&M to D R&B1&M to E R&D&E to B1 R&D&E to M R&M&D&E to B2 R&B1&M&D&E to B2 Average
Visual (Face frames) ResNet18 52.34 52.88 51.38 51.56 51.42 50.48 51.25 51.25 51.66 50.39 52.27
Visual (Face frames) ResNet18+GRU 52.11 54.04 53.25 53.43 51.04 52.22 53.26 52.18 52.36 54.32 53.96
Visual (Face frames) ResNet18+SVM 52.66 54.65 52.89 53.75 51.02 51.46 52.35 52.21 53.07 53.32 54.18
Visual (AU+Gaze) MLP 51.25 52.02 53.46 52.33 51.75 51.21 52.15 50.52 51.32 50.88 52.75
Visual (Affect) MLP 51.62 52.72 53.86 54.33 51.02 52.38 51.77 52.63 51.56 51.74 53.28
Visual (AU+Gaze+Affect) MLP 52.02 54.32 52.64 54.31 51.56 53.22 52.45 53.25 52.36 52.41 53.88
Audio (Mel spectrogram) ResNet18 50.21 51.62 51.32 52.46 50.28 50.11 51.24 51.36 50.86 52.46 52.02
Audio (Mel spectrogram) ResNet18+SVM 52.32 51.85 52.03 51.24 50.38 51.26 52.37 52.44 51.69 51.76 52.45
Audio (Waveform) Wave2Vec 52.94 53.26 52.68 53.06 52.66 52.46 52.10 52.21 51.62 52.47 52.94
@@
Modality & Inputs Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 R&M to D R&M to E R&B1 to D R&B1 to E
Visual (Face frames) ResNet18 52.00 53.47 56.44 50.00 59.81 41.58 55.45 54.28 53.22 53.27 52.88
Visual (Face frames) ResNet18+GRU 54.46 41.58 66.34 50.62 51.40 56.44 60.40 54.69 52.59 53.64 53.48
Visual (Face frames) ResNet18+SVM 55.25 51.36 62.17 52.33 51.28 52.54 53.31 52.42 53.88 54.21 52.31
Visual (AU+Gaze) MLP 51.08 43.56 55.45 53.75 57.01 53.47 54.46 51.52 52.43 53.66 51.85
Visual (Affect) MLP 55.69 57.43 57.43 51.56 52.34 49.50 61.39 53.11 53.00 54.28 52.14
Visual (AU+Gaze+Affect) MLP 50.15 56.44 58.42 50.00 57.94 60.40 63.37 54.31 53.26 53.65 53.15
Audio (Mel spectrogram) ResNet18 52.31 50.50 58.42 49.38 53.27 56.44 59.41 51.06 52.31 51.45 52.06
Audio (Mel spectrogram) ResNet18+SVM 51.36 50.67 52.31 50.42 52.57 56.38 58.01 51.88 52.04 52.35 52.38
Audio (Waveform) Wave2Vec 56.00 47.52 44.55 53.12 67.29 57.43 58.42 52.06 52.46 53.06 53.22
Modality & Inputs Method D&E to B1 D&E to B2 D&E to M D&E to R R&B1&M to D R&B1&M to E R&D&E to B1 R&D&E to M R&M&D&E to B2 R&B1&M&D&E to B2 Average
Visual (Face frames) ResNet18 52.14 53.02 52.69 51.49 56.33 52.46 52.72 52.04 51.48 51.31 52.77
Visual (Face frames) ResNet18+GRU 52.68 53.06 55.38 55.64 55.58 56.03 55.42 52.35 53.92 52.65 54.21
Visual (Face frames) ResNet18+SVM 53.03 55.25 55.06 54.23 56.63 58.02 55.37 53.24 52.85 54.31 54.24
Visual (AU+Gaze) MLP 51.94 52.31 52.58 54.62 55.31 53.27 52.48 51.56 52.47 51.78 52.69
Visual (Affect) MLP 52.15 54.36 55.40 53.19 52.27 55.43 52.52 51.23 50.82 51.25 53.64
Visual (AU+Gaze+Affect) MLP 52.39 53.48 54.28 53.62 52.25 53.56 53.06 52.18 51.59 52.46 54.28
Audio (Mel spectrogram) ResNet18 54.24 55.23 52.63 55.82 53.22 55.21 52.89 51.28 53.97 56.22 53.68
Audio (Mel spectrogram) ResNet18+SVM 52.26 54.98 53.48 55.28 52.01 53.83 53.09 52.35 55.22 51.09 53.05
Audio (Waveform) Wave2Vec 54.99 55.93 53.29 55.83 54.25 53.29 53.26 53.25 55.42 56.99 54.36
@@ 4.3 Cross-domain Testing with Unimodal Features
Intra-Domain Testing. Before performing cross-domain testing, we show the results of intra-domain testing on Real-life Trial (R), Bag of Lies (B1), MU3D (M), DOLOS (D), and MDPE (E) in Table II. It is important to note that we did not use the Box of Lies (B2) dataset for training due to concerns about its quality. Specifically, we performed a subject-independent five-fold cross-testing and reported the average accuracies. Among unimodal inputs, mel spectrograms with ResNet18 achieve the highest accuracy (75.78% on R), highlighting the strength of audio cues, particularly in structured settings like courtroom videos. Visual face frames also perform well with higher average results compared with audio features. However, facial landmarks are less reliable and show consistently low accuracy, which may due to the sparse geometric information and its sensitivity to pose, occlusion, and lighting; thus, we exclude them from the following cross-domain tasks. Multimodal fusion consistently improves performance, with the proposed Atten-Mixer achieving the best results across several datasets (e.g., 73.85% on D, 58.17% on E). Performance variation across datasets suggests that data quality, recording conditions, and content type significantly impact model effectiveness. Overall, the intra-testing results indicate the benefit of combining rich audio and visual features with advanced fusion methods for improved deception detection performances.
Single-to-Single Domain. Specifically, the models were trained on one dataset (one domain) and tested on the other dataset (another domain). As shown in Tabel III, overall, visual-based models exhibit strong generalization compared to audio-only models, particularly those leveraging rich visual descriptors (AU, gaze, affect) or deep networks like ResNet18+KNN and FFCSN. Notably, the best average performance across all transfer scenarios is achieved by FFCSN (54.50%), closely followed by visual-only model ResNet18+SVM (54.44%) and ResNet18+KNN (54.26%), highlighting the effectiveness of visual temporal modeling. Although audio contributes valuable cues, it struggles more with domain shifts compared to visual modalities. Additionally, handcrafted prosodic/acoustic features perform worse overall (e.g., MLP: 50.80% avg), compared with other types of audio features.
Multi-to-Single Domain. Table IV presents the domain generalization results under three sampling strategies: Domain-Simultaneous, Domain-Alternating, and Domain-by-Domain. Each strategy evaluates 21 domain generalization tasks using a consistent set of nine input features and model configurations. The final row in each strategy block reports the average accuracy across all 21 tasks, serving as an overall performance indicator for that sampling strategy. It is worth noting that these 21 tasks do not represent the full combination of the six datasets, but they cover most scenarios and are sufficiently comprehensive for evaluation. Our results indicate that the sampling strategy, feature selection, and dataset composition all play critical roles in building generalizable deception detection systems. We offer several key observations and discussions below:
(1) Among the three strategies, Domain-by-Domain achieves the best overall average accuracy (average across all the results) of 53.66%, followed by Domain-Alternating (53.06%) and Domain-Simultaneous (52.30%). This suggests that training on individual domains sequentially allows models to better abstract transferable patterns, likely due to reduced complexities between domains during training. In contrast, domain-alternating and domain-simultaneous keep switching between source domains, appearing to induce higher variability and lower overall accuracies.
(2) Across all strategies, models trained on visual features generally outperform those trained on audio features. This is shown in the average results from each block. For domain-simultaneous training, the best performance (54.60%) is achieved using visual behavioral features (AU+Gaze+Affect) with an MLP classifier. For domain-alternating, the best average performance is achieved by visual (face frames) trained with ResNet18+SVM. For domain-by-domain, although the best performance is achieved by audio (Waveform), its visual performances outperform audio ones, highlighting the robustness of visual cues over audio modalities.
(3) Regarding different features, overall, visual and behavioral features perform better than audio features. However, the performance using different strategies varies. For domain-simultaneous, we can observe that behavioral features (AUs, gaze, and affective cues) are more predictive than face frames or audio features. Their lower dimensionality and interpretability allow effective learning even on limited training data. For domain-alternating, face frames is better. For domain-by-domain, audio features achieve the best performance. Wave2Vec generally outperforms handcrafted features like Mel spectrograms or acoustic-prosodic embeddings, highlighting the importance of using pretrained self-supervised representations.
(4) Regarding dataset effects, one observation is that training on clean, lab-collected datasets (such as Bag of Lies and MU3D) tends to generalize better to real-world or noisier datasets (like Real-life Trial or Box of Lies). For example, under domain-simultaneous sampling, training on B1&M and testing on R yields the best result of 69.16%. This performance is consistent with the domain-by-domain strategy, where B1&M to R reaches 67.29%. Conversely, when training noisy or lower-quality datasets (e.g., R, E), performance often degrades, suggesting that domain complexity and data quality significantly affect generalization.